# Tree-Guided Diffusion Planner

### v1

@@ 3.3 Challenges in Guided Planning
Guided planning has primarily been evaluated on simple tasks where expert demonstrations are available or the guide function is convex and differentiable. As a result, prior works have focused on improving the sampling quality of the pretrained diffusion model, which was the main bottleneck in test-time guidance in these simplified settings Lee et al. (2023); Chen et al. (2024b). Despite the importance of improving trajectory sampling quality, there has been limited investigation into how guidance algorithms themselves adapt to increasing task complexity and respond to various forms of test-time objectives. In this section, we first study a fundamental challenge in guided planning: the exploitation-exploration trade-off, and then discuss the limitations of naïve gradient guidance, a standard test-time planning approach.
@@
Exploration-Exploitation Trade-off.   Test-time guided planning employs two distinct score functions: score estimate from the pretrained diffusion model and a user-defined guide score. Not only to generate a feasible trajectory but also to maximize its fitness, the agent is encouraged to balance the exploitation of the pretrained model and the exploration of novel trajectories. Gradient-based guidance typically requires selecting a guidance strength $\alpha$ to balance adherence to the guide signal and trajectory fidelity. However, $\alpha$ is highly task-dependent, and exhaustive tuning across tasks introduces significant overhead during evaluation. For example, Fig. 1 illustrates several gold-picking tasks within a fixed maze map. The optimal value of $\alpha$ varies across tasks, depending on how far the gold location deviates from unconditional navigation trajectory, which favors the shortest path (more details in Supplementary material).
@@
In-distribution Trajectory Preference.   Diffusion models are capable of generating compositional behaviors Du et al. (2020); Janner et al. (2022), but often get stuck in local optima due to insufficient exploration of the trajectory space. Most previous works on guided planning did not adequately address the exploration-exploitation trade-off, as they typically assume convex (or concave) guidance with a single optimal trajectory that globally maximizes the reward Liang et al. (2023); Lu et al. (2025). However, real-world tasks often involve multiple objectives with different priorities, where the agent must explore the trajectory space to avoid sub-optimality. Pretrained diffusion planners tend to favor generating in-distribution trajectories that align with previously seen data, rather than discovering novel compositional solutions Chen et al. (2024b). Consequently, gradient-based guidance algorithms do not effectively address this dilemma, as they often prioritize local optimal trajectories within the learned distribution (see Fig. 2-a).
@@
Non-differentiable Rule.  Gradient-based guidance algorithms face significant challenges in planning tasks with non-differentiable constraints. Since diffusion planners are trained on trajectories with fixed start and end states, they struggle to generate feasible paths that incorporate additional intermediate goals. For example, as described in Fig. 2 -b, navigation trajectories conditioned to test-time intermediate goal often result in suboptimal (left) or infeasible (right). This introduces a non-differentiable constraint from the planner’s perspective, as the requirement to pass through a specific state imposes a discrete structural condition not reflected in the training distribution. Similar challenges arise in other domains of diffusion-based generation, such as enforcing chord progression in music generation Huang et al. (2024) or satisfying chemical rules in molecule generation Shen et al. (2025)—both of which introduce hard constraints that are difficult to optimize using standard gradient-based guidance.
@@ 4 Method
Appendix˜AB outline the overall TDP pipeline and present the full algorithms. We detail the core modules of TDP—state decomposition (Sec. 4.1), parent branching (Sec. 4.2), and sub-tree expansion (Sec. 4.3).
@@ 4.1 State Decomposition
For instance, the KUKA robot arm environment provides state vectors containing multiple features such as robot joint angles and block positions. Since TDP operates as a zero-shot planner, it does not have prior knowledge of the state category for each feature of the state vector. Given a test-time block stacking task with a distance-based guide function, TDP autonomously categorizes each feature value in the state vector using Algorithm˜1. It evaluates whether the gradient of the guide function with respect to the ith feature (i.e., $\frac{\partial\mathcal{J}}{\partial\boldsymbol{s}_{i}}$) is zero or non-zero. If non-zero, the ith feature is classified as an observation state; if zero, it is classified as a control state. Consequently, features related to robot physics are detected as control states, while block position (xy) features are detected as observation states, because the block-stacking guide function is only affected by the position of the blocks.
@@ 4.2 Parent Branching
In the first phase of our bi-level planning framework, control states are applied to fixed-potential particle guidance (PG) Corso et al. (2024) to explore diverse control trajectories. PG promotes diversity among generated samples within a batch. For instance, when moving a block to a target position, multiple control trajectories can accomplish this task. Standard gradient-based approaches often suffer from in-distribution bias and limited exploration, constraining trajectory diversity. In contrast, TDP enhances exploration through this procedure, generating what we term parent trajectories.
@@
Specifically, fixed-potential PG is implemented with the gradient of a radial basis function (RBF), denoted $\nabla\Phi$, which can be directly computed with all pairwise distances between control trajectories within a batch. Unlike conventional gradient guidance methods that pull samples toward high-reward regions, PG introduces repulsive forces that push samples apart in the data space. This leads to a broad coverage of dynamically feasible trajectories independent of task objectives. Although fixed-potential PG incurs some inference overhead from computing kernel values between all trajectory pairs, it is significantly more sample-efficient than learned-potential variants Corso et al. (2024), while remaining training-free and modular. A single denoising step for parent branching is denoted as
@@
$\boldsymbol{\mu}^{i-1}_{\text{control}}\leftarrow\boldsymbol{\mu}^{i}_{\text{control}}+\alpha_{p}\Sigma^{i}\nabla\Phi(\boldsymbol{\mu}^{i}_{\text{control}}),\quad\boldsymbol{\mu}^{i-1}_{\text{obs}}\leftarrow\boldsymbol{\mu}^{i}_{\text{obs}}+\alpha_{g}\Sigma^{i}\nabla\mathcal{J}(\boldsymbol{\mu}^{i}_{\text{obs}})$ (3)
@@
where $\boldsymbol{\mu}^{i}_{\text{control}}$ and $\boldsymbol{\mu}^{i}_{\text{obs}}$ denote the control and observation components of the predicted mean of the denoising trajectory at timestep $i$. Particle guidance $\nabla\Phi(\boldsymbol{\mu}_{\text{control}})$ introduces repulsive updates among control states to diversify denoising paths, while gradient guidance $\nabla\mathcal{J}(\boldsymbol{\mu}_{\text{obs}})$ steers observation states toward task-relevant regions defined by the guide function. This enables a wide exploration in the control state space, which helps to discover diverse observation state configurations and exposes the planner to richer gradient signals from the guide function $\mathcal{J}$.
@@
Notably, TDP formulates a single joint conditional distribution with an integrated guidance term. As shown in Equation˜2, the overall guidance distribution is defined as $h(\boldsymbol{\tau}_{i})=h_{\text{gg}}(\boldsymbol{\tau}_{i})\cdot h_{\text{pg}}(\boldsymbol{\tau}_{i})$ where $h_{\text{gg}}(\cdot)$ denotes the gradient guidance component, and $h_{\text{pg}}(\cdot)$ denotes the particle guidance component. Since both components jointly condition the same pretrained reverse denoising process, the perturbed reverse denoising process is approximated as $\tilde{p}(\boldsymbol{\tau}_{i-1}|\boldsymbol{\tau}_{i})\approx\mathcal{N}(\boldsymbol{\tau}_{i-1};\mu^{i}+\alpha_{\text{TDP{}}}\Sigma^{i}g_{\text{TDP{}}},\Sigma^{i})$ where $g_{\text{TDP{}}}=\nabla_{\tau}\log\left(h_{\text{gg}}(\boldsymbol{\tau}_{i})\cdot h_{\text{pg}}(\boldsymbol{\tau}_{i})\right)=\nabla_{\tau}\log h_{\text{gg}}(\boldsymbol{\tau}_{i})+\nabla_{\tau}\log h_{\text{pg}}(\boldsymbol{\tau}_{i})=g_{\text{gg}}+g_{\text{pg}}$. Therefore,
@@
denotes the integrated guidance term from Equation˜3, where $\boldsymbol{\mu}^{i-1}=[\boldsymbol{\mu}^{i-1}_{\text{control}},\boldsymbol{\mu}^{i-1}_{\text{obs}}]$ and $g_{\text{TDP{}}}=g_{\text{gg}}+g_{\text{pg}}=\frac{1}{\alpha_{\text{TDP{}}}}\left(\alpha_{p}\nabla\Phi(\boldsymbol{\mu}^{i}_{\text{control}})+\alpha_{g}\nabla\mathcal{J}(\boldsymbol{\mu}^{i}_{\text{obs}})\right)$. TDP employs a single integrated guidance term, which is the sum of the gradient guidance part ($g_{\text{gg}}$) and the particle guidance part ($g_{\text{pg}}$).
@@ 4.3 Sub-Tree Expansion
$\boldsymbol{\tau}_{\text{child}}^{N_{f}}$ $\sim q_{N_{f}}(\boldsymbol{\tau}_{\text{parent}},\boldsymbol{C})\quad\text{where}\;\boldsymbol{C}=\{\boldsymbol{s}_{k}\}_{k=0}^{b}\;\text{and}\;b\sim Unif(0,T_{\text{pred}})$ (5)
@@
We investigate the role of our bi-level sampling framework in handling multi-reward structures, as illustrated in Figure˜3. We characterize problems containing both local and global optima, demonstrating how bi-level trajectory generation avoids local optima. Consider trajectory data in a learned subspace, where the pretrained diffusion planner maps Gaussian noise back to data space via the reverse denoising process. Given a test-time guide function defined as the sum of Gaussian components (as defined in Proposition 1), the following statements hold:
@@
Initialization problem in gradient guidance with diffusion planner. Assume that the trajectory data $X\in\mathbb{R}^{H\times D}$ follows the Assumption 1 in Guo et al. (2024), and given guide $\mathcal{J}(X)=\mathcal{J}_{1}(X)+\mathcal{J}_{2}(X)$ where $\mathcal{J}_{1}(X)=\exp\left(-\frac{1}{2\sigma_{1}^{2}}\|X-Av_{1}\|^{2}\right)$, $\mathcal{J}_{2}(X)=\exp\left(-\frac{1}{2\sigma_{2}^{2}}\|X-w_{\perp}\|^{2}\right)$, $w_{\perp}\perp span(A)$, $v_{1}$ is the first eigenvector of $A^{\top}A$, $\sigma_{1}<\sigma_{2}$, and $\mathbb{E}_{n\sim\mathcal{N}(0,I)}[\mathcal{J}_{1}(n)]\ll\mathbb{E}_{n\sim\mathcal{N}(0,I)}[\mathcal{J}_{2}(n)]$.
@@
The proof is provided in Appendix C. Guided sampling initialized from standard Gaussian noise can be steered away from the subspace, converging to off-subspace local optima Guo et al. (2024), as illustrated in Proposition 1-a. These samples may locally optimize the task objective but lie outside the learned data manifold, resulting in unrealistic or infeasible trajectories. This occurs because standard gradient guidance modifies the pretrained denoising process without preserving the underlying data structure, causing the sampling to drift toward regions that satisfy the guidance objective but violate the training distribution constraints. In practice, many test-time planning problems exhibit off-subspace optima—for example, in maze navigation, unseen obstacles can block the learned shortest path, making pretrained trajectory preferences suboptimal. In contrast, unconditional samples naturally remain close to the learned data subspace since they follow the original training distribution. When guided sampling is initialized from these on-subspace points, it can leverage both data structure and gradient information to converge to the global optimum, as demonstrated in Proposition 1-b.
TDP’s bi-level planning framework enables flexible exploration through structural search over trajectories. During parent branching, gradient guidance can be optionally applied to steer the observation states, depending on task characteristics. Unconditional PG supports broad exploration of the data space, effective for discovering diverse solutions in under-specified or multi-goal settings. In contrast, conditional PG directs exploration toward regions aligned with the guide function, useful for convex objectives where a single high-quality solution is desired (e.g., PnP tasks in Sec. 5.3). Gradient guidance in parent branching provides a prior for sub-tree expansion but may also limit exploration breadth in trajectory space. Each strategy offers complementary benefits: unconditional PG promotes diversity without task-specific priors, while conditional PG enables targeted search.
@@ 5 Experiments
While existing diffusion planners (e.g., MCTD Yoon et al. (2025), Hierarchical Diffuser Chen et al. (2024b)) excel on standard offline benchmarks, they suffer from zero-shot scenarios as discussed in Sec. 2. Consequently, we focus comparisons on recent zero-shot planning approaches, particularly Trajectory Aggregation Tree (TAT) Feng et al. (2024), Monte-Carlo sampling Lu et al. (2025), and stochastic sampling Wang et al. (2025). We design our benchmarks to challenge planners with unseen test-time objectives, contrasting with offline benchmarks that only test learned dynamics aligned with the training distribution. For completeness, we provide supplementary comparisons with sequential approaches (i.e., MCTD Yoon et al. (2025), Diffusion-Forcing Chen et al. (2024a)) on standard maze benchmarks in Appendix J. Notably, despite being designed specifically for zero-shot planning, TDP still surpasses these sequential approaches on standard benchmarks.

### v2

@@ 3.3 Challenges in Guided Planning
Guided planning has primarily been evaluated on simple tasks where expert demonstrations are available or the guide function is convex and differentiable. As a result, prior work has focused on improving the sampling quality of the pretrained diffusion model, which was the main bottleneck for test-time guidance in these simplified settings [chen2024simplehierarchicalplanningdiffusion, lee2023refining]. Despite the importance of improving trajectory sampling quality, there has been limited investigation into how guidance algorithms themselves adapt to increasing task complexity and respond to various forms of test-time objectives. In this section, we first study a fundamental challenge in guided planning: the exploitation-exploration trade-off, and then discuss the limitations of naïve gradient guidance, a standard test-time planning approach.
@@
Exploration-Exploitation Trade-off.   Test-time guided planning employs two distinct score functions: score estimate from the pretrained diffusion model and a user-defined guide score. Not only to generate a feasible trajectory but also to maximize its fitness, the agent is encouraged to balance the exploitation of the pretrained model and the exploration of novel trajectories. Gradient-based guidance typically requires selecting a guidance strength $\alpha$ to balance adherence to the guide signal and trajectory fidelity. However, $\alpha$ is highly task-dependent, and exhaustive tuning across tasks introduces significant overhead during evaluation. For example, Fig. 1 illustrates several gold-picking tasks within a fixed maze map. The optimal value of $\alpha$ varies across tasks, depending on how far the gold location deviates from the unconditional navigation trajectory, which favors the shortest path (see the supplement for more details).
@@
In-distribution Trajectory Preference.   Diffusion models are capable of generating compositional behaviors [janner2022planningdiffusionflexiblebehavior, du2020compositionalvisualgenerationinference], but often get stuck in local optima due to insufficient exploration of the trajectory space. Most previous works on guided planning did not adequately address the exploration-exploitation trade-off, as they typically assume convex (or concave) guidance with a single optimal trajectory that globally maximizes the reward [liang2023adaptdiffuserdiffusionmodelsadaptive, lu2025what]. However, real-world tasks often involve multiple objectives with different priorities, where the agent must explore the trajectory space to avoid sub-optimality. Pretrained diffusion planners tend to favor generating in-distribution trajectories that align with previously seen data, rather than discovering novel compositional solutions [chen2024simplehierarchicalplanningdiffusion]. Consequently, gradient-based guidance algorithms do not effectively address this dilemma, as they often prioritize local optimal trajectories within the learned distribution (see Fig. 2-a).
@@
Non-differentiable Rule.  Gradient-based guidance algorithms face significant challenges in planning tasks with non-differentiable constraints. Since diffusion planners are trained on trajectories with fixed start and end states, they struggle to generate feasible paths that incorporate additional intermediate goals. For example, as described in Fig. 2 -b, navigation trajectories conditioned to test-time intermediate goal often result in suboptimal (left) or infeasible (right). This introduces a non-differentiable constraint from the planner’s perspective, as the requirement to pass through a specific state imposes a discrete structural condition not reflected in the training distribution. Similar challenges arise in other domains of diffusion-based generation, such as enforcing chord progression in music generation [huang2024symbolicmusicgenerationnondifferentiable] or satisfying chemical rules in molecule generation [shen2025chemistryinspired]. Both introduce hard constraints that are difficult to optimize with standard gradient-based methods.
@@ 4 Method
Appendix˜AB outline the overall TDP pipeline and present the full algorithms. We detail the core modules of TDP: state decomposition (Sec. 4.1), parent branching (Sec. 4.2), and sub-tree expansion (Sec. 4.3).
@@ 4.1 State Decomposition
For instance, the KUKA robot arm environment provides state vectors containing multiple features such as robot joint angles and block positions. Since TDP operates as a zero-shot planner, it does not have prior knowledge of the state category for each feature of the state vector. Given a test-time block stacking task with a distance-based guide function, TDP autonomously categorizes each feature value in the state vector using Algorithm˜1. It evaluates whether the gradient of the guide function with respect to the ith feature (i.e., $\frac{\partial\mathcal{J}}{\partial\boldsymbol{s}_{i}}$) is zero or non-zero. If non-zero, the ith feature is classified as an observation state; if zero, it is classified as a control state. Consequently, features related to robot physics are detected as control states. In contrast, block position (xy) features are detected as observation states, because the block-stacking guide function is only affected by the block positions.
@@ 4.2 Parent Branching
In the first phase of our bi-level planning framework, control states are applied to fixed-potential particle guidance (PG) [corso2024particle] to explore diverse control trajectories. PG promotes diversity among generated samples within a batch. For instance, when moving a block to a target position, multiple control trajectories can accomplish this task. Standard gradient-based approaches often suffer from in-distribution bias and limited exploration, constraining trajectory diversity. In contrast, TDP enhances exploration through this procedure, generating what we term parent trajectories.
@@
Specifically, fixed-potential PG is implemented using the gradient of a radial basis function (RBF), denoted $\nabla\Phi$, which can be computed directly from all pairwise distances between control trajectories within a batch. Unlike conventional gradient guidance methods that pull samples toward high-reward regions, PG introduces repulsive forces that push samples apart in the data space. This leads to a broad coverage of dynamically feasible trajectories independent of task objectives. Although fixed-potential PG incurs some inference overhead from computing kernel values between all trajectory pairs, it is significantly more sample-efficient than learned-potential variants [corso2024particle], while remaining training-free and modular. A single denoising step for parent branching is denoted as:
@@
$\boldsymbol{\mu}^{i-1}_{\text{control}}\leftarrow\boldsymbol{\mu}^{i}_{\text{control}}+\alpha_{p}\Sigma^{i}\nabla\Phi(\boldsymbol{\mu}^{i}_{\text{control}}),\quad\boldsymbol{\mu}^{i-1}_{\text{obs}}\leftarrow\boldsymbol{\mu}^{i}_{\text{obs}}+\alpha_{g}\Sigma^{i}\nabla\mathcal{J}(\boldsymbol{\mu}^{i}_{\text{obs}}),$ (3)
@@
where $\boldsymbol{\mu}^{i}_{\text{control}}$ and $\boldsymbol{\mu}^{i}_{\text{obs}}$ denote the control and observation components of the predicted mean of the denoising trajectory at timestep $i$. Particle guidance $\nabla\Phi(\boldsymbol{\mu}_{\text{control}})$ introduces repulsive updates among control states to diversify denoising paths. In contrast, gradient guidance $\nabla\mathcal{J}(\boldsymbol{\mu}_{\text{obs}})$ steers observation states toward task-relevant regions defined by the guide function. This enables a wide exploration in the control state space, which helps to discover diverse observation state configurations and exposes the planner to richer gradient signals from the guide function $\mathcal{J}$.
@@
Notably, TDP formulates a single joint conditional distribution with an integrated guidance term. As shown in Eq. 2, the overall guidance distribution is defined as $h(\boldsymbol{\tau}_{i})=h_{\text{gg}}(\boldsymbol{\tau}_{i})\cdot h_{\text{pg}}(\boldsymbol{\tau}_{i})$ where $h_{\text{gg}}(\cdot)$ denotes the gradient guidance component, and $h_{\text{pg}}(\cdot)$ denotes the particle guidance component. Since both components jointly condition the same pretrained reverse denoising process, the perturbed reverse denoising process is approximated as $\tilde{p}(\boldsymbol{\tau}_{i-1}|\boldsymbol{\tau}_{i})\approx\mathcal{N}(\boldsymbol{\tau}_{i-1};\mu^{i}+\alpha_{\text{TDP{}}}\Sigma^{i}g_{\text{TDP{}}},\Sigma^{i})$ where $g_{\text{TDP{}}}=\nabla_{\tau}\log\left(h_{\text{gg}}(\boldsymbol{\tau}_{i})\cdot h_{\text{pg}}(\boldsymbol{\tau}_{i})\right)=\nabla_{\tau}\log h_{\text{gg}}(\boldsymbol{\tau}_{i})+\nabla_{\tau}\log h_{\text{pg}}(\boldsymbol{\tau}_{i})=g_{\text{gg}}+g_{\text{pg}}$. Therefore,
@@
denotes the integrated guidance term from Eq. 3, where $\boldsymbol{\mu}^{i-1}=[\boldsymbol{\mu}^{i-1}_{\text{control}},\boldsymbol{\mu}^{i-1}_{\text{obs}}]$ and $g_{\text{TDP{}}}=g_{\text{gg}}+g_{\text{pg}}=\frac{1}{\alpha_{\text{TDP{}}}}\left(\alpha_{p}\nabla\Phi(\boldsymbol{\mu}^{i}_{\text{control}})+\alpha_{g}\nabla\mathcal{J}(\boldsymbol{\mu}^{i}_{\text{obs}})\right)$. TDP employs a single integrated guidance term, which is the sum of the gradient guidance part ($g_{\text{gg}}$) and the particle guidance part ($g_{\text{pg}}$).
@@ 4.3 Sub-Tree Expansion
$\boldsymbol{\tau}_{\text{child}}^{N_{f}}$ $\sim q_{N_{f}}(\boldsymbol{\tau}_{\text{parent}},\boldsymbol{C})\quad\text{where}\;\boldsymbol{C}=\{\boldsymbol{s}_{k}\}_{k=0}^{b}\;\text{and}\;b\sim Unif(0,T_{\text{pred}}),$ (5)
@@
We investigate the role of our bi-level sampling framework in handling multi-reward structures, as illustrated in Fig. 3. We characterize problems with both local and global optima and demonstrate that bi-level trajectory generation avoids local optima. Consider trajectory data in a learned subspace, where the pretrained diffusion planner maps Gaussian noise back to data space via the reverse denoising process. Given a test-time guide function defined as the sum of Gaussian components (as defined in Proposition 1), the following statements hold:
@@
Initialization problem in gradient guidance with diffusion planner. Assume that the trajectory data $X\in\mathbb{R}^{H\times D}$ follows the Assumption 1 in [guo2024gradientguidancediffusionmodels], and given guide $\mathcal{J}(X)=\mathcal{J}_{1}(X)+\mathcal{J}_{2}(X)$ where $\mathcal{J}_{1}(X)=\exp\left(-\frac{1}{2\sigma_{1}^{2}}\|X-Av_{1}\|^{2}\right)$, $\mathcal{J}_{2}(X)=\exp\left(-\frac{1}{2\sigma_{2}^{2}}\|X-w_{\perp}\|^{2}\right)$, $w_{\perp}\perp span(A)$, $v_{1}$ is the first eigenvector of $A^{\top}A$, $\sigma_{1}<\sigma_{2}$, and $\mathbb{E}_{n\sim\mathcal{N}(0,I)}[\mathcal{J}_{1}(n)]\ll\mathbb{E}_{n\sim\mathcal{N}(0,I)}[\mathcal{J}_{2}(n)]$.
@@
The proof is provided in Appendix C. Guided sampling initialized from standard Gaussian noise can be steered away from the subspace, converging to off-subspace local optima [guo2024gradientguidancediffusionmodels], as illustrated in Proposition 1-a. These samples may locally optimize the task objective but lie outside the learned data manifold, resulting in unrealistic or infeasible trajectories. This occurs because standard gradient guidance modifies the pretrained denoising process without preserving the underlying data structure, causing the sampling to drift toward regions that satisfy the guidance objective but violate the training distribution constraints. In practice, many test-time planning problems exhibit off-subspace optima—for example, in maze navigation, unseen obstacles can block the learned shortest path, making pretrained trajectory preferences suboptimal. In contrast, unconditional samples naturally remain close to the learned data subspace since they follow the original training distribution. When guided sampling is initialized from these on-subspace points, it can leverage both data structure and gradient information to converge to the global optimum, as demonstrated in Proposition 1-b.
TDP’s bi-level planning framework enables flexible exploration through structural search over trajectories. During parent branching, gradient guidance can be optionally applied to steer the observation states, depending on task characteristics. Unconditional PG supports broad exploration of the data space, effective for discovering diverse solutions in under-specified or multi-goal settings. In contrast, conditional PG directs exploration toward regions aligned with the guide function, useful for convex objectives where a single high-quality solution is desired (e.g., PnP tasks in Sec. 5.3). Gradient guidance in parent branching provides a prior for sub-tree expansion but may also limit the breadth of exploration in trajectory space. Each strategy offers complementary benefits: unconditional PG promotes diversity without task-specific priors, while conditional PG enables targeted search.
@@ 5 Experiments
While existing diffusion planners (e.g., MCTD [yoon2025montecarlotreediffusion], Hierarchical Diffuser [chen2024simplehierarchicalplanningdiffusion]) excel on standard offline benchmarks, they suffer from zero-shot scenarios as discussed in Sec. 2. Consequently, we focus comparisons on recent zero-shot planning approaches, particularly Trajectory Aggregation Tree (TAT) [feng2024resistingstochasticrisksdiffusion], Monte-Carlo sampling [lu2025what], and stochastic sampling [wang2025inferencetimepolicysteeringhuman]. We design our benchmarks to challenge planners with unseen test-time objectives, in contrast to offline benchmarks that only test learned dynamics aligned with the training distribution. For completeness, we provide supplementary comparisons with sequential approaches (i.e., MCTD [yoon2025montecarlotreediffusion], Diffusion-Forcing [chen2024diffusionforcingnexttokenprediction]) on standard maze benchmarks in Appendix J. Notably, despite being designed specifically for zero-shot planning, TDP still surpasses these sequential approaches on standard benchmarks.